If the Lagrangian L is thus given by L(x, ) =  show that Lagrange’s

 equation of motion becomes


1
Expert's answer
2021-05-25T16:16:08-0400
"\frac{d}{dt}\frac{\partial L}{\partial v}=\frac{\partial L}{\partial x}"